OpenStudy (anonymous):

-1/2 divided by 3/4?

OpenStudy (anonymous):

= -1/2 multiplied by 4/3

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by its reciprocal which is what telltoamit is showing you

OpenStudy (anonymous):

-4/6

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

-4/6 = -2/3

OpenStudy (anonymous):

thanks.